# Multiple solutions to a semilinear elliptic equation with a sharp change of sign in the nonlinearity

### v1

@@ Appendix A A generic property
[…]
@@
equipped with the norm $\|\cdot\|_{m}$. Let $\mathfrak{B}_{\rho}:=\left\{\theta\in\mathfrak{C}^{2}\ :\ \|\theta\|_{2}\leq\rho\right\}$ be the ball of radius $\rho$ in $\mathfrak{C}^{2}$ centered at 00, and set
@@
There exists $\rho_{0}>0$ such that, if $\theta\in\mathfrak{B}_{\rho}$, $\rho\in(0,\rho_{0})$, and 00 is a regular value of the function $x\mapsto F(x,\theta)$, then $\psi_{\Omega_{\theta}}$ is a Morse function.
@@
By definition, 00 is a regular value of $x\mapsto F(x,\theta)$ if and only if $D_{x}F(x,\theta)$ is an isomorphism for every $x$ such that $F(x,\theta)=0$. Now, if $\rho$ is small enough, then $I+D\theta(x)$ is a linear isomorphism. Therefore, $F(x,\theta)=0$ if and only if $\nabla\psi_{\Omega_{\theta}}(x+\theta(x))=0$, i.e., if and only if $x+\theta(x)$ is a critical point of $\psi_{\Omega_{\theta}}$. Similarly, if $D_{x}F(x,\theta)$ is an isomorphism, then $x+\theta(x)$ is nondegenerate. ∎
@@
$\{\theta\in\mathfrak{B}_{\rho}:0\text{ is a regular value of the function }x\mapsto F(x,\theta)\}$
@@
We show that, for small enough $\rho$, the function $F$ defined in (A.1) satisfies the conditions $i)$, $ii)$, and $iii)$ of Theorem A.3 with $\Omega=:U\subset X:=\mathbb{R}^{n}$, $\mathfrak{B}_{\rho}=:V\subset Y:=\mathfrak{C}^{2}$, and $Z=\mathbb{R}^{n}$.
$i):$ This assumption is trivially satisfied because every $\mathcal{C}^{1}$-map $\Omega\to\mathbb{R}^{n}$ is Fredholm of index 00.
@@
and $\overline{\mathfrak{B}}_{r}$ is the closed ball of radius $r$ in $\mathfrak{C}^{2}$ centered at 00. Clearly, $F^{-1}(0)=\cup_{j\in\mathbb{N}}A_{j}$. Next, we show that the restriction $\pi|_{A_{j}}$ to $A_{j}$ of the projection $\pi:\mathbb{R}^{n}\times\mathfrak{C}^{2}\to\mathfrak{C}^{2}$ is a proper map for every $j\in\mathbb{N}$. Let $\mathfrak{K}$ be a compact subset of $\mathfrak{C}^{2}$ and $(x_{k},\theta_{k})$ be a sequence in $A_{j}\cap\pi^{-1}(\mathfrak{K})$. Then there is a subsequence such that $x_{k}\to x$ in $\Omega_{j}^{-}$ and $\theta_{k}\to\theta$ in $\mathfrak{K}$ (because $\mathfrak{K}$ is compact and $(\theta_{k})\subset\mathfrak{K}$). This shows that $A_{j}\cap\pi^{-1}(\mathfrak{K})$ is compact, i.e., $\pi|_{A_{j}}$ is proper.
$ii):$ Let $(x,\theta)\in\Omega\times\mathfrak{B}_{\rho}$. The change of variables $z=y+\theta(y)$ yields
@@
Fix $(x,\theta)\in\Omega\times\mathfrak{B}_{\rho}$ such that $F(x,\theta)=0$. To prove condition $ii)$, it suffices to show that
@@
where $\mathbb{H}_{\nu_{\zeta}}:=\{z\in\mathbb{R}^{n}:\langle z,\nu_{\zeta}\rangle>0\}$ (here we used the smoothness of the domain). Now, $(I+D\theta(\zeta))^{-1}=I+A(\zeta)$, where the entries of $A(\zeta)$ tend to 00 as $\rho\to 0$. Therefore,
@@
All we need to do now is to show that there exist $\zeta_{1},\ldots,\zeta_{n}\in\partial\Omega$ such that $x-\zeta_{1}+\theta(x)-\theta(\zeta_{1}),\ldots,x-\zeta_{n}+\theta(x)-\theta(\zeta_{n})$ are linearly independent. To this end, let $\{e_{1},\ldots,e_{n}\}$ be the canonical basis in $\mathbb{R}^{n}$. Then, there is $t_{i}>0$ such that $\widetilde{\zeta}_{i}=x+\theta(x)+t_{i}e_{i}\in\partial\Omega_{\theta}$ for every $i=1,\ldots,n$. Let $\zeta_{i}\in\partial\Omega$ be such that $\widetilde{\zeta}_{i}=\zeta_{i}+\theta(\zeta_{i})$. Then, $x-\zeta_{i}+\theta(x)-\theta(\zeta_{i})=-t_{i}e_{i}$, so these vectors are linearly independent and, thus, $ii)$ is satisfied. The claim now follows from Theorem A.3. ∎
@@ Appendix B An invertibility result
$|f_{0}(U_{1}-U_{2})-f_{0}(U_{1})+f_{0}(U_{2})|_{\frac{(p+1)(1+\gamma)}{p}}=\begin{cases}O(\delta^{n-2})&\text{ if }3\leq n<6,\\ O(\delta^{n-2})&\text{ if }n=6\text{ and }\gamma<0,\\ O(\delta^{4}|\log\delta|^{\frac{2}{3}})&\text{ if }n=6\text{ and }\gamma=0,\\ O(\delta^{\frac{n+2}{2}})&\text{ if }n>6,\end{cases}$
@@
[…]
@@
$\int_{\omega_{+}}\left||W|^{p-1}W-U_{1}^{p}+U_{2}^{p}\right|^{\frac{2n}{n+2}}$ =$=\begin{cases}O(\delta^{\frac{2n(n-2)}{n+2}})&\text{ if }3\leq n<6,\\ O(\delta^{6}|\log\delta|)&\text{ if }n=6,\\ O(\delta^{n(1+\gamma)})&\text{ if }n>6\end{cases}$
@@
$\int_{\omega_{+}}\left||W|^{p-1}W-U_{1}^{p}+U_{2}^{p}\right|^{\frac{(p+1)(1+\gamma)}{p}}$ =$=\begin{cases}O(\delta^{\frac{2n(n-2)(1+\gamma)}{n+2}})&\text{ if}3\leq n\leq 6,\\ O(\delta^{n(1+\gamma)})&\text{ if }n>6.\end{cases}$

### v2

@@ Appendix A A generic property
$\psi_{\Omega}(x):=\int_{\mathbb{R}^{n}\setminus\Omega}\frac{1}{|x-y|^{2n}}dy,\qquad x\in\Omega,$
@@
equipped with the norm $\|\cdot\|_{m}$. Let $\mathfrak{B}_{\rho}:=\left\{\theta\in\mathfrak{C}^{2}\ :\ \|\theta\|_{2}\leq\rho\right\}$ be the ball of radius $\rho$ in $\mathfrak{C}^{2}$ centered at 0, and set
@@
There exists $\rho_{0}>0$ such that, if $\theta\in\mathfrak{B}_{\rho}$, $\rho\in(0,\rho_{0})$, and 0 is a regular value of the function $x\mapsto F(x,\theta)$, then $\psi_{\Omega_{\theta}}$ is a Morse function.
@@
By definition, 0 is a regular value of $x\mapsto F(x,\theta)$ if and only if $D_{x}F(x,\theta)$ is an isomorphism for every $x$ such that $F(x,\theta)=0$. Now, if $\rho$ is small enough, then $I+D\theta(x)$ is a linear isomorphism. Therefore, $F(x,\theta)=0$ if and only if $\nabla\psi_{\Omega_{\theta}}(x+\theta(x))=0$, i.e., if and only if $x+\theta(x)$ is a critical point of $\psi_{\Omega_{\theta}}$. Similarly, if $D_{x}F(x,\theta)$ is an isomorphism, then $x+\theta(x)$ is nondegenerate. ∎
@@
[…]
@@
We show that, for small enough $\rho$, the function $F$ defined in (A.1) satisfies the conditions $i)$, ii)ii), and $iii)$ of Theorem A.3 with $\Omega=:U\subset X:=\mathbb{R}^{n}$, $\mathfrak{B}_{\rho}=:V\subset Y:=\mathfrak{C}^{2}$, and $Z=\mathbb{R}^{n}$.
$i):$ This assumption is trivially satisfied because every $\mathcal{C}^{1}$-map $\Omega\to\mathbb{R}^{n}$ is Fredholm of index 0.
@@
and $\overline{\mathfrak{B}}_{r}$ is the closed ball of radius $r$ in $\mathfrak{C}^{2}$ centered at 0. Clearly, $F^{-1}(0)=\cup_{j\in\mathbb{N}}A_{j}$. Next, we show that the restriction $\pi|_{A_{j}}$ to $A_{j}$ of the projection $\pi:\mathbb{R}^{n}\times\mathfrak{C}^{2}\to\mathfrak{C}^{2}$ is a proper map for every $j\in\mathbb{N}$. Let $\mathfrak{K}$ be a compact subset of $\mathfrak{C}^{2}$ and $(x_{k},\theta_{k})$ be a sequence in $A_{j}\cap\pi^{-1}(\mathfrak{K})$. Then there is a subsequence such that $x_{k}\to x$ in $\Omega_{j}^{-}$ and $\theta_{k}\to\theta$ in $\mathfrak{K}$ (because $\mathfrak{K}$ is compact and $(\theta_{k})\subset\mathfrak{K}$). This shows that $A_{j}\cap\pi^{-1}(\mathfrak{K})$ is compact, i.e., $\pi|_{A_{j}}$ is proper.
ii):ii): Let $(x,\theta)\in\Omega\times\mathfrak{B}_{\rho}$. The change of variables $z=y+\theta(y)$ yields
@@
Fix $(x,\theta)\in\Omega\times\mathfrak{B}_{\rho}$ such that $F(x,\theta)=0$. To prove condition ii)ii), it suffices to show that
@@
where $\mathbb{H}_{\nu_{\zeta}}:=\{z\in\mathbb{R}^{n}:\langle z,\nu_{\zeta}\rangle>0\}$ (here we used the smoothness of the domain). Now, $(I+D\theta(\zeta))^{-1}=I+A(\zeta)$, where the entries of $A(\zeta)$ tend to 0 as $\rho\to 0$. Therefore,
@@
All we need to do now is to show that there exist $\zeta_{1},\ldots,\zeta_{n}\in\partial\Omega$ such that $x-\zeta_{1}+\theta(x)-\theta(\zeta_{1}),\ldots,x-\zeta_{n}+\theta(x)-\theta(\zeta_{n})$ are linearly independent. To this end, let $\{e_{1},\ldots,e_{n}\}$ be the canonical basis in $\mathbb{R}^{n}$. Then, there is $t_{i}>0$ such that $\widetilde{\zeta}_{i}=x+\theta(x)+t_{i}e_{i}\in\partial\Omega_{\theta}$ for every $i=1,\ldots,n$. Let $\zeta_{i}\in\partial\Omega$ be such that $\widetilde{\zeta}_{i}=\zeta_{i}+\theta(\zeta_{i})$. Then, $x-\zeta_{i}+\theta(x)-\theta(\zeta_{i})=-t_{i}e_{i}$, so these vectors are linearly independent and, thus, ii)ii) is satisfied. The claim now follows from Theorem A.3. ∎
@@ Appendix B An invertibility result
[…]
@@
$\int_{\omega_{+}}U_{2}^{(p+1)(1+\gamma)}=\int_{\omega_{+}}\left(\frac{\alpha_{n}^{p+1}\delta^{n}}{(\delta^{2}+|x-\xi_{2}|^{2})^{n}}\right)^{1+\gamma}\,dx\leq C|\omega_{+}|\delta^{n(1+\gamma)}=O(\delta^{n(1+\gamma)}).$ (B.3)
@@
$\int_{\omega_{+}}\left||W|^{p-1}W-U_{1}^{p}+U_{2}^{p}\right|^{\frac{2n}{n+2}}$ $=\begin{cases}O(\delta^{\frac{2n(n-2)}{n+2}})&\text{ if }3\leq n<6,\\ O(\delta^{6}|\log\delta|)&\text{ if }n=6,\\ O(\delta^{n(1+\gamma)})&\text{ if }n>6\end{cases}$
@@
$\int_{\omega_{+}}\left||W|^{p-1}W-U_{1}^{p}+U_{2}^{p}\right|^{\frac{(p+1)(1+\gamma)}{p}}$ $=\begin{cases}O(\delta^{\frac{2n(n-2)(1+\gamma)}{n+2}})&\text{ if }3\leq n\leq 6,\\ O(\delta^{n(1+\gamma)})&\text{ if }n>6.\end{cases}$
@@
Acknowledgments
We thank the anonymous referees for their careful reading and valuable comments, which helped improve the quality of the paper. M. Clapp and A. Saldaña thank La Sapienza Università di Roma for the kind hospitality during the preparation of this work.